verify cauchy's mean value theorem for the function f(x) = x and g(x) = sinx in [0,π/2].

1
Expert's answer
2021-12-30T12:58:36-0500

The functions "f(x)=x" and "g(x)=\sin x" are continuous on the interval "[0,\pi/2]," differentiable on "(0,\pi/2)," and "g'(x)=\cos x≠0" for all "x\in(0,\pi/2)."

Then there is a point "x=c" in this interval such that


"\dfrac{f(\pi/2)-f(0)}{g(\pi/2)-g(0)}=\dfrac{f'(c)}{g'(c)}"

"f(\pi/2)=\pi/2, f(0)=0"

"g(\pi/2)=\sin(\pi/2)=1, g(0)=\sin(0)=0"

"f'(x)=1, f'(c)=1"

"g'(x)=\cos x, g'(c)=\cos c"

"\dfrac{\pi/2-0}{1-0}=\dfrac{1}{\cos c}"

"\cos c=2/\pi, c\in(0,\pi/2)"

"c=\cos^{-1}(2/\pi)"

It is evident that this number "c=\cos^{-1}(2/\pi)" lies in the interval "(0,\pi/2),"

 i.e. satisfies the Cauchy theorem.